In bolt factory , machine A ,B ,and C manufacturer 25, 35 and 40

Percent of the total output , 5, 4 and 2 percent , respectively are

defective bolts.A bolts is selected at random and found to be defective.

What is the probability that the bolt came from machine A , B and C ?
1
Expert's answer
2020-07-21T17:15:58-0400

here, "P(A)= \\frac {25}{100}\n\u200b", "P(B)= \\frac {35}{100}\n\u200b", "P(C)= \\frac {40}{100}\n\u200b"


"P(D\/A)= \\frac {5}{100}\n\u200b", "P(D\/B)= \\frac {4}{100}\n\u200b", "P(D\/C)= \\frac {2}{100}\n\u200b"

where D denotes defective bolts


"P(D)=P(A).P( D\/A )+P(B).P(D\/B )+P(C).P( D\/C)"


"P(D)=\\frac {25}{100}.\\frac {5}{100}+\\frac {35}{100}\n\u200b\t\\frac {4}{100} +\\frac {40}{100}\n\u200b\t\\frac {2}{100}" "= 0.0345"


"P(A/D)=\frac {P(A).P( D/A)} {P(A).P( D/A)+P(B).P(D/B )+P(C).P( D/C)}"


"P(A/D)=\frac {\frac{25}{100}. \frac {5}{100} } {0.0345}=\frac {25}{69}"



"P(B/D)=\frac {P(B).P( D/B)} {P(A).P( D/A)+P(B).P(D/B )+P(C).P( D/C)}"


"P(B/D)=\frac {\frac{35}{100}. \frac {4}{100} } {0.0345}=\frac {28}{69}"


"P(C/D)=\frac {P(C).P( D/C )} {P(A).P( D/A )+P(B).P(D/B)+P(C).P( D/C)}"


"P(C/D)=\frac {\frac{40}{100}. \frac {2 }{100} } {0.0345}=\frac {16}{69}"